OpenStudy (anonymous):

can someone help me to understand 3 1/8 x 1 1/5.

OpenStudy (texaschic101):

Turn them into improper fractions and just multiply. Do you know how to turn mixed numbers int improper fractions ?

OpenStudy (anonymous):

ya

OpenStudy (texaschic101):

So what is 3 1/8 as an improper fraction ?

OpenStudy (anonymous):

so it would be like 25/8 x 6/5 = ?

OpenStudy (zehanz):

Remember 3 1/8= 3+1/8. so convert the 3 wholes to 8ths and the 1 to 5ths. Then you cann write 3 1/8 as a fraction and 1 1/5 as a fraction. Now you have two fractions. Multiply them by this rule:\[\frac{ a }{ b }\cdot \frac{ c }{ d }=\frac{ ac }{ bd }\]

OpenStudy (texaschic101):

yes...exactly

OpenStudy (anonymous):

ok

OpenStudy (zehanz):

Also, simplify!

OpenStudy (texaschic101):

Do you understand ?

OpenStudy (anonymous):

yes thank you

OpenStudy (texaschic101):

your quite welcome

OpenStudy (texaschic101):

thank you for the medal :)

OpenStudy (texaschic101):

thank you Zehanz :)

OpenStudy (zehanz):

@texaschic101: YW!